# LeapFrog: The Rowhammer Instruction Skip Attack

### v1

@@ 5.1 Offline Memory Profiling
Finding Contiguous Memory
In this work, we employed DRAMA [38] to ascertain contiguous memory regions. DRAMA provides a direct and efficient method for reverse engineering the DRAM layout, facilitating the identification of contiguous memory. DRAMA exploits the physical structure of DRAM through a timing side channel to reveal addressing information, enabling us to determine contiguous memory regions and bank affiliations. This is crucial for Rowhammer, as the attack requires the targeted rows to be in adjacent physical locations within the same DRAM bank.
To optimize memory allocation, we define a structure that maps the DRAM and virtual addresses of each page, then allocate a large buffer to populate this structure with the corresponding addresses. By implementing a two-phase sorting process—initially sorting by DRAM bank, followed by sorting rows within each bank—we efficiently identify sequences of physically contiguous rows within the same bank. This streamlined approach eliminates redundancy and precisely targets contiguous memory regions, enhancing the overall memory management strategy.
After sorting, we traverse the array to find the longest sequence of contiguous rows within the same bank. This is achieved by comparing the DRAM row values, looking for sequences that are the same or increment by one. The result is a segment of memory, identified as a continuous chunk, which is critical for mounting a successful Rowhammer attack.
Executing the Rowhammer Bit Flip in a Multi-Sided Context Despite modern mitigation techniques against Rowhammer like Target Row Refresh (TRR), we are still able to induce flips in DDR4 memory by using a multi-sided [18] approach.
In the final phase of our attack, the task is to induce bit flips in the target memory location. This step marks the culmination of the profiling and memory manipulation processes. The challenge lies in the fact that while we can ascertain the occurrence of bit flips in a given row (a row that we deem "flippy"), pinpointing the exact memory bits affected after the attack is not straightforward. This is due to the inherent nature of Rowhammer, where the attacker does not possess direct control over the specific memory areas being altered.
However, the success of the attack is often evident through observable changes in the process’ state. For instance, a successful execution might manifest as an unauthorized bypass of security measures, or broken encryption output. This indirect outcome serves as a confirmation of the attack’s effectiveness. We further expand on this in §7.
@@ 6 Locating the PC in the Stack
To flip bits in the PC value with Address Space Layout Randomization (ASLR) enabled, the page that contains the PC value needs to be placed into the page with the bit that will flip during the Rowhammer attack. To do this, we use a method similar to that proposed in [40] where we deallocate a series of pages from the attacker process, launch the victim process, and experimentally determine some probability that the target data (in this case the PC) lands in the target location (the row with the flippy bits). We term the deallocation of pages "baiting" in this paper.
@@
The dynamic nature of the PC under ASLR implemented in the Linux kernel, necessitates a novel approach that involves identifying invariant values within the stack that serve as reliable fingerprints. These fingerprints are used to determine the PC’s offset relative to these constants, thereby facilitating the estimation of the required number of bait pages for effective targeting.
@@ 6.1 Fingerprinting the Stack
As the PC’s address and value fluctuate with each process execution due to ASLR, our strategy leverages the relative stability of certain stack values and correlating an offset from those values. We first profile with ASLR disabled, knowing the target PC value in the stack from an assembly dump with GDB. We then determine an offset from the fingerprint as seen in Figure 4. Then with ASLR enabled, even with the PC value changing, the fingerprint remains identifiable and the offset from the fingerprint remains constant. The outcome is a refined understanding of the number of bait pages required to strategically position the PC, thus enhancing the precision of our Rowhammer attack in an ASLR-enabled environment. Fingerprinting only needs to be done once and is machine-independent.
@@ 7 Automatic Detection of LeapFrog Gadgets with MFS
Based on how the LeapFrog gadgets occur in the binary described in §5, we developed a custom tool we call MFS (Multidimensional Fault Simulator) that relies on dynamic binary instrumentation and analysis. Since the attack happens on program counters and registers, which are invisible to high-level code, such as C/C++, it is not possible to do a static analysis of the source code. We put together a set of rules that enables us to collect, filter, and pinpoint the potential LeapFrog gadgets.
@@
First, MFS collects the instruction traces, specifically, the address of instructions executed, for different inputs. For the purpose of detecting the gadgets that cause security exploits, MFS chooses critical input pairs that cause differences in the program’s control flow. Such inputs can be correct/incorrect private key pairs or passphrases for authentication programs. Together with the instruction addresses, we collect the execution time of each function executed. Since the return addresses of the functions with larger execution times will stay in the memory for a longer duration, they are potentially more viable targets.
@@
MFS then computes the difference between two instruction traces to find the instruction addresses that are executed with correct input(s) but not executed with incorrect input(s). Note that this is an optional step to reduce the complexity of the following steps, and it comes with a cost of false negatives. Moreover, depending on the program and type of exploit, it may not always be possible to get multiple different traces; see §8.1. Alternatively, the whole instruction trace can be considered instead of only the difference.
@@
MFS then looks for address pairs that hold the following conditions:
$d_{H}(\textrm{addr}_{\text{exec}}^{i},\textsf{addr}_{\text{return}}^{j})=1$ (1)
where $addr_{\text{exec}}^{i}$ is the address of the $i^{th}$ instruction that is executed, $addr_{\text{return}}^{j}$ is the return addresses of the $j^{th}$ call instruction, and $d_{H}$ is the Hamming distance between two addresses. $i$ and $j$ are bounded by the number of all instructions executed ($n$) and the number of call instructions executed ($m$), respectively. Although this operation has $O(m^{n})$ complexity, it can be implemented with bitwise xor and can be parallelized using multiple processor cores. The condition given in Equation (1) is determined by the Rowhammer fault model. Since multiple-bit flips on a precise target are much rarer and harder to control, MFS assumes we can only flip a single bit. Yet, the method is generic enough to cover other potential fault models, such as optical fault injection, where multiple-bit flips are more likely [9]. This step generates a list of pairs of addresses in the following format: $\{<addr_{src}^{k},addr_{dest}^{k}>\}$ where $addr_{src}^{k}$ is the $k^{th}$ instruction address that MFS targets in the binary’s execution with the input that we want to affect the control flow of, such as an incorrect private key, and $addr_{dest}^{k}$ is the corrupted instruction address after fault injection.
@@
For each address pair we get from the list generated in the previous step, MFS starts a simulation session. MFS executes the binary again with the incorrect input and simulate a bit flip on the instruction address $addr_{src}$ to make it $addr_{dest}$. It is possible that certain instructions are executed multiple times in a single execution. To correctly cover that case in our fault model, we keep a counter variable for a specific instruction that increments every time the binary executes the same instruction. In a single execution of the original binary, if an instruction is executed NN times, we attempt the fault simulation $N+1$ times, until we no longer see the same instruction in the trace.
@@
After the bit flip simulation, MFS continues the execution of the binary without further faults and observe the new behavior. The analysis of the new behavior is not a trivial task. There are several options where we can observe changes compared to the original execution. For instance, we can observe changes in the total number of executed instructions, the number of instructions that match with the correct input execution trace, the return code of the program, outputs to standard streams, ports that are accessed, functions calls, authentication result, etc. The choice of observable depends on the program under test. In this work, MFS uses the return codes, standard outputs/errors, and authentications on different case studies.
@@
Once MFS has a list of PC values that potentially result in authentication or bypass with a single-bit flip, it then evaluates if they are practical to attack from a timing perspective. In some cases, a single bit flip will result in the desired behavior in a process but the attack window of time is too short to effectively attack the target. Additionally, the attack window needs to be long enough to allow for noise in the system - as processes will often take a variable amount of time to execute and get to the vulnerable area in the code where the PC value is shelved in the stack. MFS uses process degradation to increase the viability of LeapFrog gadgets, as slowing down a process artificially increases the attack window time. Note that this step is system-specific and it can be affected by the current processor/memory load. Although it is necessary to find viable targets in the list for an end-to-end attack, it does not guarantee that the other targets are not viable in different system configurations, or in different systems.
MFS starts the victim process and then immediately stops it with a SIGSTOP signal and it checks if the PC value is currently in the stack of the process. If not, the process is killed and restarted, and stopped after a slightly longer period of time, in a process we call time sweeping. The challenge is sending a SIGSTOP with the highest timing precision possible. Different implementations of signals will yield different timing resolutions. For example, Python has a signal library that can be used to generate signals similar to a bash script, but there is considerable delay and imprecision in the time it takes to send a signal.
@@ 7.1 Tool Implementation
We used Intel’s dynamic binary instrumentation framework, Pin, which allow for process analysis without altering its core behavior  [32] to implement 1 and 4 of MFS. Using Pin also makes it possible to find LeapFrog gadgets in binaries that do not have a source code since it does not require recompiling. In the context of MFS, Pin’s capabilities are harnessed to monitor the execution trace of a binary. This integration allows for a thorough analysis of potential LeapFrog gadgets by observing how changes in PC values influence program behavior. For each executed instruction, our tool outputs the virtual address of the instruction and disassembly of the machine code. If the instruction is a call instruction it also outputs the return address of the call, which is usually the PC value that is pushed onto the stack before executing the called routine. For every write to STDOUT and STDERR, the tool forwards a copy of the buffer to a text file for further analysis. To avoid the effect of overhead caused by instruction-based instrumentation, function timings are collected in a separate session on every function entry and exit.
@@
The bit flip simulation part of MFS ( 4) is done using Pin which takes the address pairs and simulates every fault independently. The faults on PC values are implemented as direct jumps to the corrupted addresses by adding jmp $addr_{dest}$ after function returns. Since we add a direct jump to the target address by injecting a line of assembly with the the Pin tool, it is functionally equivalent to corrupting the PC value in memory.
@@
filters the simulation results depending on the program and targeted exploit type. For different types of exploits, we filter by return code (§8.3), value in STDOUT (§8.1), and authentication result (§8.2).
@@
In order to determine if a PC value is in the stack, we start the victim process as a non-root user on a sibling core to a core that we are attacking with SMC to degrade the performance. For example, in our tests we isolated cores 6 and 14 and triggered SMC events on core 14 while running the victim process on core 6. Once we have the process identification number (PID) of the process and send a SIGSTOP, we use the Linux process interface to check the stack for the PC value. We do this by looking at the /proc/[pid]/maps file to determine which offsets in the victim process’s address space contain the stack, and then we read from /proc/[pid]/mem at the offsets determined by /proc/[pid]/maps to find the PC values. The tool will generate a dictionary of stack addresses/values for the victim process that we can search through.
@@ 8 Experiments
Experiment Setup The experiments are conducted on a system with Ubuntu 22.04.2 LTS with 6.2.0-37-generic Linux kernel installed. The system uses an Intel Core i9-9900K CPU with a Coffee Lake microarchitecture. We used a dynamic clock frequency rather than a static clock frequency to improve the practicality of the attack. End-to-end attack experiments are done on a single DIMM Corsair DDR4 DRAM chip with part number CMU64GX4M4C3200C16 and 16GB capacity. DRAM row refresh period is kept at 64ms, which is the default value in most systems. In all the experiments, we used 100s simulation timeout, since the fault simulations rarely cause infinite loops. We empirically observe that using the Python signals library, the target process could complete 34M cycles before the attacker can stop it, with a standard deviation of 2.7M cycles. Alternatively using a bash script, the victim process can only complete 18M cycles before it is stopped, with a standard deviation of 0.3M cycles. There is an order of magnitude difference in precision stopping a process with bash vs with Python.
Target Size #Inst.exec $d_{HD}$ # Candidates

### v2

@@ 5.1 Offline Memory Profiling
Finding Contiguous Memory Virtual to physical address mappings are stored in pagemap file in Linux OSs and it requires root privileges to access these translations. Using the SPOILER tool [28], we can reliably leak the information about the first 8 bits in physical addresses after the page offset bits. This allows us to find contiguous memory chunks in physical address space. In Figure 3, the page numbers with the peaks in the y-axis are contiguous pages’ physical address space.
Figure 3: Timing peaks on virtual addresses detected by SPOILER [28] attack. Virtual addresses on the peaks are contiguous on physical address space.
Finding Memory In the Same Banks To find physical memory pages within the same bank, we employ techniques first described in [46]. DRAM is structured in multiple banks that are physically isolated from each other, so while SPOILER can give an attacker physically contiguous chunks, the memory is distributed across multiple banks. This is a problem because hammering rows that are not adjacent within the same bank will not result in bit flips.
We use the row conflict side channel to find co-located memory address accesses within the same bank. We take measurements by iterative reading between two addresses, and if the reading results in high latency (assuming we are not hitting cache), it means DRAM is clearing the row buffer and the addresses are located in the same bank.
Alternatively, if we repeatedly read from two addresses and they are in different banks, the values will be loaded into their respective row buffers for their respective banks, and the reading time will have a lower latency.
Executing the Rowhammer Bit Flip in a Many-Sided Context Despite modern mitigation techniques against Rowhammer like Target Row Refresh (TRR), we are still able to induce flips in DDR4 memory by using a many-sided [21] approach.
In the final phase of our attack, the task is to induce bit flips in the target memory location. This step marks the culmination of the profiling and memory manipulation processes. The challenge lies in the fact that while we can ascertain the occurrence of bit flips in a given row (a row that we deem “flippy”), pinpointing the exact memory bits affected after the attack is not straightforward. This is due to the inherent nature of Rowhammer, where the attacker does not possess direct control over the specific memory areas being altered.
However, the success of the attack is often evident through observable changes in the process’ state. For instance, a successful execution might manifest as an unauthorized bypass of security measures, or broken encryption output. This indirect outcome serves as a confirmation of the attack’s effectiveness. We further expand on this in section 7.
@@ 6 Locating the PC in the Stack
To flip bits in the PC value with Address Space Layout Randomization (ASLR) enabled, the page that contains the PC value needs to be placed into the page with the bit that will flip during the Rowhammer attack. To do this, we use a method similar to that proposed in [50] where we deallocate a series of pages from the attacker process, launch the victim process, and experimentally determine some probability that the target data (in this case the PC) lands in the target location (the row with the flippy bits). We term the deallocation of pages “baiting” in this paper.
@@
The dynamic nature of the PC under ASLR implemented in the Linux kernel necessitates a novel approach that involves identifying invariant values within the stack that serve as reliable fingerprints. These fingerprints are used to determine the PC’s offset relative to these constants, thereby facilitating the estimation of the required number of bait pages for effective targeting.
@@ 6.1 Fingerprinting the Stack
As the PC’s address and value fluctuate with each process execution due to ASLR, our strategy leverages the relative stability of certain stack values and correlates an offset from those values. We first profile with ASLR disabled, knowing the target PC value in the stack from an assembly dump with GDB. We then determine an offset from the fingerprint as seen in Figure 4. Then with ASLR enabled, even with the PC value changing, the fingerprint remains identifiable and the offset from the fingerprint remains constant. The outcome is a refined understanding of the number of bait pages required to strategically position the PC, thus enhancing the precision of our Rowhammer attack in an ASLR-enabled environment. Fingerprinting only needs to be done once and is machine-independent.
@@ 7 Automatic Detection of LeapFrog Gadgets with MFS
Based on how the LeapFrog gadgets occur in the binary described in section 5, we develop a custom tool we call MFS (Multidimensional Fault Simulator) that relies on dynamic binary instrumentation and analysis. Since the attack happens on program counters and registers, which are invisible to high-level code, such as C/C++, it is not possible to do a static analysis of the source code. We put together a set of rules that enables us to collect, filter, and pinpoint the potential LeapFrog gadgets. The overall design is shown in Figure 6.
@@
First, MFS collects the instruction traces, specifically, the address of instructions executed, for different inputs. To detect the gadgets that cause security exploits, MFS chooses critical input pairs that cause differences in the program’s control flow. Such inputs can be correct/incorrect private key pairs or passphrases for authentication programs. Together with the instruction addresses, we collect the execution time of each function executed. Since the return addresses of the functions with larger execution times will stay in the memory for a longer duration, they are potentially more viable targets.
@@
MFS then computes the difference between two instruction traces to find the instruction addresses that are executed with correct input(s) but not executed with incorrect input(s). Note that this is an optional step to reduce the complexity of the following steps, and it comes with a cost of false negatives. Moreover, depending on the program and type of exploit, it may not always be possible to get multiple different traces; see section 8.2. Alternatively, the whole instruction trace can be considered instead of only the difference.
@@
Regarding the choice of the fault model on PC values, the probability analysis in previous work [59] is relevant. Given a sequence of bit offsets ${b_{0},b_{1},...,b_{k+l-1}}$ within a memory page and assuming a defective memory cell can flip solely in one direction, the conditional probability of identifying a compatible target page $t$ amongst $N$ susceptible pages can be expressed as:
$p\big{(}t|\{b_{n_{0\rightarrow 1}}\}\in\{0\rightarrow 1\},\{b_{n_{1\rightarrow 0}}\}\in\{1\rightarrow 0\}\big{)}=\\ 1-\bigg{(}1-\prod_{i=0}^{k-1}{\dfrac{n_{0\rightarrow 1}-i}{S-i}}\times\prod_{j=0}^{l-1}{\dfrac{n_{1\rightarrow 0}-j}{S-k-j}}\bigg{)}^{N},$ (1)
where $n_{0\rightarrow 1}$ and $n_{1\rightarrow 0}$ represent the average counts of error-prone cells on a page that can be flipped from 00 to $1$ and from $1$ to 00 respectively, $k$ and $l$ denote the counts of bit locations needing flips from 00 to $1$ and $1$ to 00 respectively, and “$S$” signifies the total bit count per page.
First, we calculate the probability of finding a target page t for each $N$ value and three different $k+l$ values. Note that $k+l$ is the number of bit offsets within a page. Figure 7 shows that for 1 bit per page, 2200 pages are enough to achieve 99.99% accuracy for a DDR4 DRAM with an average of 100 flips per page. For 2 and 3 bits per page, the same number of pages gives 2% and 0.006% probability, respectively.
Figure 7: Probability of finding a page among N pages for different k+l values. k+l states the number of targeted bit offsets in a page.
MFS looks for address pairs that hold the following conditions:
$d_{HD}(\textrm{addr}_{\text{exec}}^{i},\textsf{addr}_{\text{return}}^{j})=1$ (2)
where $addr_{\text{exec}}^{i}$ is the address of the $i^{th}$ instruction that is executed, $addr_{\text{return}}^{j}$ is the return addresses of the $j^{th}$ call instruction, and $d_{HD}$ is the Hamming distance between two addresses. $i$ and $j$ are bounded by the number of all instructions executed ($n$) and the number of call instructions executed ($m$), respectively. Although this operation has $O(m^{n})$ complexity, it can be implemented with bitwise xor and can be parallelized using multiple processor cores. The condition given in Equation (2) is determined by the Rowhammer fault model. Given that finding a suitable memory page in memory is only realistic with single-bit flip fault models, MFS assumes we can only flip a single bit. Yet, the method is generic enough to cover other potential fault models, such as optical fault injection or electromagnetic fault injection, where multiple-bit flips are more likely [10]. This step generates a list of pairs of addresses in the following format: $\{<addr_{src}^{k},addr_{dest}^{k}>\}$ where $addr_{src}^{k}$ is the $k^{th}$ instruction address that MFS targets in the binary’s execution with the input that we want to affect the control flow of, such as an incorrect private key, and $addr_{dest}^{k}$ is the corrupted instruction address after fault injection.
@@
For each address pair we get from the list generated in the previous step, MFS starts a simulation session. MFS executes the binary again with the incorrect input and simulates a bit flip on the instruction address $addr_{src}$ to make it $addr_{dest}$. Certain instructions may be executed multiple times in a single execution. To correctly cover that case in our fault model, we keep a counter variable for a specific instruction that increments every time the binary executes the same instruction. In a single execution of the original binary, if an instruction is executed $N$ times, we attempt the fault simulation $N+1$ times, until we no longer see the same instruction in the trace.
@@
After the bit flip simulation, MFS continues the execution of the binary without further faults and observes the new behavior. The analysis of the new behavior is not a trivial task. There are several options where we can observe changes compared to the original execution. For instance, we can observe changes in the total number of executed instructions, the number of instructions that match with the correct input execution trace, the return code of the program, outputs to standard streams, ports that are accessed, function calls, authentication results, etc. The choice of observable depends on the program under test. In this work, MFS uses the return codes, standard outputs/errors, and authentications on different case studies.
@@
Once MFS has a list of PC values that potentially result in misauthentication or bypass with a single-bit flip, it then evaluates if they are practical to attack from a timing perspective. In some cases, a single bit flip will result in the desired behavior in a process but the attack window of time is too short to effectively attack the target. Additionally, the attack window needs to be long enough to allow for noise in the system - as processes will often take a variable amount of time to execute and get to the vulnerable area in the code where the PC value is shelved in the stack. MFS uses process degradation to increase the viability of LeapFrog gadgets, as slowing down a process artificially increases the attack window time. Note that this step is system-specific and it can be affected by the current processor/memory load. Although it is necessary to find viable targets in the list for an end-to-end attack, it does not guarantee that the other targets are not viable in different system configurations, or different systems.
MFS starts the victim process and then immediately stops it with a SIGSTOP signal and it checks if the PC value is currently in the stack of the process. If not, the process is killed and restarted, and stopped after a slightly longer period, in a process we call time sweeping. The challenge is sending a SIGSTOP with the highest timing precision possible. Different implementations of signals will yield different timing resolutions. For example, Python has a signal library that can be used to generate signals similar to a bash script, but there is considerable delay and imprecision in the time it takes to send a signal.
@@ 7.1 Tool Implementation
We used Intel’s dynamic binary instrumentation framework, Pin [38], which allows for process analysis without altering its core behavior to implement 1 and 4 of MFS. Using Pin also makes it possible to find LeapFrog gadgets in binaries that do not have a source code since it does not require recompiling. In the context of MFS, Pin’s capabilities are harnessed to monitor the execution trace of a binary. This integration allows for a thorough analysis of potential LeapFrog gadgets by observing how changes in PC values influence program behavior. For each executed instruction, our tool outputs the virtual address of the instruction and disassembly of the machine code. If the instruction is a call instruction, it also outputs the return address of the call. The return address of the call is usually the PC value that is pushed onto the stack before executing the called routine. For every write to STDOUT and STDERR, the tool forwards a copy of the buffer to a text file for further analysis. To avoid the effect of overhead caused by instruction-based instrumentation, function timings are collected in a separate session on every function entry and exit.
@@
The bit flip simulation part of MFS ( 4) is done using Pin which takes the address pairs and simulates every fault independently. The faults on PC values are implemented as direct jumps to the corrupted addresses by adding jmp $addr_{dest}$ after function returns. Since we add a direct jump to the target address by injecting a line of assembly with the Pin tool, it is functionally equivalent to corrupting the PC value in memory.
@@
filters the simulation results depending on the program and targeted exploit type. For different types of exploits, we filter by return code (section 8.4), value in STDOUT (section 8.2).
@@
To determine if a PC value is in the stack, we start the victim process as a non-root user on a sibling core to a core that we are attacking with SMC to degrade the performance. For example, in our tests we isolated cores 6 and 14 and triggered SMC events on core 14 while running the victim process on core 6. Once we have the process identification number (PID) of the process and send a SIGSTOP, we use the Linux process interface to check the stack for the PC value. We do this by looking at the /proc/[pid]/maps file to determine which offsets in the victim process’s address space contain the stack, and then we read from /proc/[pid]/mem at the offsets determined by /proc/[pid]/maps to find the PC values. The tool will generate a dictionary of stack addresses/values for the victim process that we can search through.
@@ 8 Experiments
Experiment Setup The experiments are conducted on a system with Ubuntu 22.04.2 LTS with 6.2.0-37-generic Linux kernel installed. The system uses an Intel Core i9-9900K CPU with a Coffee Lake microarchitecture. We used a dynamic clock frequency instead of a static clock frequency to improve the practicality of the attack. End-to-end attack experiments are performed on a single DIMM Corsair DDR4 DRAM chip with part number CMU64GX4M4C3200C16 and 16GB capacity. DRAM row refresh period is kept at 64ms, which is the default value in most systems. In all the experiments, we used 100s simulation timeout, since the fault simulations rarely cause infinite loops. We empirically observe that using the Python signals library, the target process could complete 34M cycles before the attacker can stop it, with a standard deviation of 2.7M cycles. Alternatively, using a bash script, the victim process can only complete 18M cycles before it is stopped, with a standard deviation of 0.3M cycles. There is an order-of-magnitude difference in precision stopping a process with bash vs. with Python.
8.1 ML Misclassification
In this section, we investigate the potential implications of instruction skipping in the machine learning domain, specifically for decision tree algorithms. A decision tree is an ML model used to make predictions based on a series of binary choices, effectively splitting data into increasingly specific groups. It starts with a single node, which branches into possible outcomes based on the features of the data. Each branch represents a decision pathway, and each node in the pathway represents a test on a specific attribute. This process continues until a leaf node is reached, which provides the predicted outcome. They are widely used in various applications, from financial forecasting [37] to medical diagnosis [54] due to their interpretability, and efficiency for a variety of tasks such as classification, and feature importance ranking. We choose a decision tree for proof of concept yet instruction skipping attacks can be effective in every kind of model implementation.
Classification algorithms may be vulnerable to the LeapFrog attack under the threat model that an attacker is co-located on the server with the victim process running the model, and the attacker would like to force a particular output. If the attacker faults the victim process program counter and forces a jump in the code, the result may be a misclassification or a forced classification of a particular output. This attack is different from other Rowhammer attacks on machine learning models [59] because for this attack we do not need to know the model weights before hand, and we consider this a gray box model.
In this experiment, we use a public implementation[44] as our target. We simulate program counter flips and observe the effects on the model output. We follow a similar procedure to previous examples, where we experiment with a hammering distance of 1, 2, and 3 and determine the number of successful LeapFrog gadgets with each of these distances. In Table I, we can see various number of LeapFrog candidate gadgets that might result in a misclassification. After simulating these gadgets, we found 23 of the 1363 potential gadgets within 1 hammer distance would result in a misclassification.
Target Size #Inst.exec $d_{HD}$ Candidates